OpenStudy (anonymous):

integral in C; In=1/(z^2+2z+4) with C={|z|=1}

OpenStudy (anonymous):

\[I_{n}= \int\limits_{C}^{}1/(z ^{2}+2z+4) with C={|z|=1}\]

OpenStudy (anonymous):

\[I _{n}=\int\limits_{C}^{}=1/(z ^{2}+2z+4)z^{n}\]

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

parametrize \(|z|=1\) as \(z=e^{it}\) \(0\leq t <2\pi\) then \[\int_0^{2\pi}f(e^{it})z'(t)dt\]

OpenStudy (anonymous):

\[\int_0^{2\pi}\frac{ie^{it}}{e^{2it}+2e^{it}+4}\]

OpenStudy (anonymous):

I tried but it equal a integral very difficult

OpenStudy (zarkon):

use residue theory

OpenStudy (anonymous):

I trying

OpenStudy (zarkon):

also...I can't fully tell what you integral is

OpenStudy (zarkon):

is it \[\oint\limits_{|z|=1}\frac{1}{z ^{2}+2z+4}dz\] ?

OpenStudy (anonymous):

miss z^n

OpenStudy (zarkon):

\[I_n=\oint\limits_{|z|=1}\frac{1}{z^n(z ^{2}+2z+4)}dz\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hmmm none of the residues are inside the unit circle though...

OpenStudy (zarkon):

the only pole you have inside \(|z|=1\) is zero. find the residue there

OpenStudy (anonymous):

?

OpenStudy (zarkon):

z=0 is a pole of order n

OpenStudy (anonymous):

oh i see i there is a \(z^n\) there let me be quiet

OpenStudy (anonymous):

ok z=0 is singular point

OpenStudy (anonymous):

yeah i am a moron integral i wrote would be zero because function has no poles inside unit circle

OpenStudy (anonymous):

thanks @satellite73 and @Zarkon I trying use residuees theory

OpenStudy (anonymous):

@Zarkon you think residues =??

OpenStudy (zarkon):

what do you think it is

OpenStudy (anonymous):

I think it is [1/(z1-z2)]*[1/z1^n]*[1/z2^n] with z1,z2 are solution of z^2+2z+4

OpenStudy (zarkon):

doesn't your formula depend in the order of z1-z2. that seems weird

OpenStudy (zarkon):

I get \[I_n=2\pi i \times \left\{\begin{matrix}0 & n\equiv 0 (\bmod3) \\ \frac{1}{2^{n+1}} & n\equiv 1(\bmod 3)\\\frac{-1}{2^{n+1}} & n\equiv 2 (\bmod3) \end{matrix}\right.\]

OpenStudy (anonymous):

you calculation \[\lim_{z \rightarrow 0}zf(z)\]

OpenStudy (zarkon):

I looked at the Laurent expansion

OpenStudy (zarkon):

if you take that limit you get 1/4

OpenStudy (zarkon):

at least for n=1

OpenStudy (anonymous):

example of the Laurent expansion? we have \[\int\limits_{C}^{}f(z)=2piiRes[f(z),z=z _{0} with Resf(z)=\lim_{z \rightarrow z _{0}}(z-z _{0})f(z)\]

OpenStudy (zarkon):

that is only for a pole of order 1

OpenStudy (zarkon):

your pole is of order n

OpenStudy (anonymous):

the pole point is 0

OpenStudy (zarkon):

yes....and it is of order n